# A radial scalar product for Kerr quasinormal modes

### v3

@@ VII.2.1 Construction
The polynomial coefficients, $\text{a}_{pkq}$, are determined by applying the form above for $\mathbf{y}_{pk}(\xi)$ to the eigen-relationship given by Eq. (73a). One then asserts that the coefficients of each power in $\xi^{q}$ must be equal on both sides of Eq. (73a). The result is a three term recursion relationship or, equivalently, a discrete eigenvalue relationship. To simplify our notation, we will henceforth drop two indices, $p$ and $k$, from $\text{a}_{pkq}$,
@@
Furthermore, if $\mu_{p}$ is given by Eq. (80), then the discrete eigenvalue relationship given by Eq. (76) defines a finite dimensional matrix eigenvalue problem151515 Note that, along with the series’ boundary conditions (Eqs. 78-79), it is the finite nature of the system of equations that justifies the eigenproblem interpretation. In contrast, non-polynomial solutions to the radial problem will not satisfy Eq. (79), meaning that e.g. a continued fraction based method, and a related boundary condition at the infinite end of the series are required Leaver (1986, 1985).  Ronveaux and Arscott (1995); DLMF (b); Cook and Zalutskiy (2014). The matrix of interest, $\hat{\mathbf{A}}_{p}$, is asymmetric and tridiagonal,
@@ VII.2.2 Orthogonality between polynomials of like order
$(\lambda_{pk}-\lambda_{pk^{\prime}})\,{\langle{\mathbf{y}_{pk^{\prime}}}\,|\,{\mathbf{y}_{pk}}\rangle}\;$ = 0.absent 0=\;0\;.= 0 . (91)
@@
Figure 5: Distribution of roots for confluent Heun polynomials, $y_{pk}(\xi)$, for three cases, all with $s=-2$, $a/M=0.7$, and $\ell=m=2$. The three cases differ in polynomial order, $p$, and the frequency, $\tilde{\omega}$, used: (left) $y_{33}$, the $n=0$ QNM frequency, $M\tilde{\omega}=0.5326-0.0808i$, was used, (center) $y_{77}$, the $n=0$ QNM frequency, $M\tilde{\omega}=0.4713-0.5843i$, was used, and (right) $y_{20,20}$, the $n=12$ QNM frequency, $M\tilde{\omega}=0.4155-2.5050i$, was used. Polynomial roots are marked by white open circles. Roots were calculated using numpy.roots Horn and Johnson (1990); Harris et al. (2020). Values of polynomial phase, $\arg(y_{pk})$, are shaded from blue to yellow, indicating values between $-\pi$ and $\pi$, respectively. Dotted horizontal and vertical lines mark the respective locations of the real and imaginary axes.
@@ VII.2.3 A scalar product expression for eigenvalues
If $y_{pk}$ are known, then Eq. (99) allows the computation of the related eigenvalues via the scalar product.
@@ VII.2.4 Numerical examples
Numerical computation of the confluent Heun polynomials may proceed by first computing $\text{C}_{2}$, $\text{C}_{3}$ and $\text{C}_{4}$. Then, for a chosen polynomial order, one would compute $\hat{\mathbf{A}}_{p}$ (Eq. 81), so that its eigenvalues and vectors may be determined. Since scalar products between polynomials is a weighted sum over monomial moments (Eq. 64), it is useful to precompute and store all of the monomial moments relevant to a given polynomial order. This allows all polynomial scalar products to be evaluated without explicit integration (Eq. V.2). From this perspective, numerical evaluation of confluent Heun polynomials may be accomplished on-the-fly via the appropriate sum over monomials (Eq. 74). Here, this procedure is used to provide select numerical examples.
@@
This effect has two causes, one technical and the other mathematical. The technical cause of this feature is simply the value of the precision parameter, dps$=64$, which means that numerical results near $10^{-64}$ are more prone to round-off error. The mathematical cause of this feature is that, like the legendre polynomials, the confluent Heun polynomials have normalization constants that generally decrease as polynomial order increases [^16] For the Legendre polynomials, the analog of $\eta_{pk}^{2}$ (see Eq. 92) is $(2p+1)^{-1}$.. Consequently the action of normalization effectively scales each polynomial by a very large number, exacerbating finite precision errors. At worst, square normalization constants might be $10^{-64}$ or smaller and would thereby be corrupted by finite precision error. In turn, this would correspond to normalization constants $10^{-32}$ or smaller, and the act of normalization would be to scale the non-normalized polynomials by a corrupted number of order $10^{32}$. This situation can be avoided by increasing numerical precision.
@@
At focus in the left panel of Fig. 7 is the distribution of normalization constants. While normalization constants are found to generally decrease with polynomial order (not shown in figure), they are also found to generally decrease as $k$ increases, and they are found to generally increase as the imaginary part $\tilde{\omega}$ increases. Branching in the distribution of normalization constants is linked to branching in the respective polynomial eigenvalues. Similarly, Eq. (95) links eigenvalue branching to the distribution of $|\langle y_{pk^{\prime}}|\xi|y_{pk}\rangle|$ seen in the right column of Fig. 6.
@@ VIII Implications for analytic solutions to Teukolsky’s radial equation
In effect, Eq. (VIII) fully defines a solution ansatz. This solution ansatz may be used to investigate whether the scalar product and confluent Heun polynomials imply that Teukolsky’s radial equation is, like the angular equation, exactly tridiagonalizable. To this end, we may attempt to represent the radial equation in the basis defined by $\{u_{p}\}_{p=0}^{\infty}$.
@@
Rewriting $\mathcal{L}{{}_{\xi}}$ in terms of $\mathcal{D}_{\xi}$ (see Eq. 101) allows |$\mathcal{L}{{}_{\xi}}\,|{u_{p}}\rangle$ to be expanded as
@@
It follows from Eq. (107c) that |$\mathcal{L}{{}_{\xi}}\,|{u_{p}}\rangle$ may be rewritten as
@@
Together, Eq. (103d), Eq. (VIII) and Eq. (113) allow $\mathcal{L}_{\xi}|{u_{p}}\rangle$ to be written as
@@
The implication of Eqs. (114-VIII) is that, if the canonical polynomials $u_{n}$ can be determined, then Teukolsky’s radial operator is exactly tridiagonalizable.
@@ Appendix A Notation
=$\mathcal{L}{{}_{r}}\;=\;$ $V_{R}(r)\;+\;\Delta(r)^{-s}\partial_{r}\left[\Delta(r)^{s+1}\partial_{r}\right]\;,$ (119)
@@ Appendix B Solving Teukolsky’s angular equation with Jacobi polynomials
Equivalently, $k_{0}$, $k_{1}$ and $k_{2}$ are the singular exponents that may be found via the method of Frobenius Courant and Hilbert. (1954); Arfken et al. (2013); Pinchover and Rubinstein (2005). In effect, $k_{0}$ and $k_{1}$ remove divergences from the potential at the regular singular points, and $k_{2}$ removes the potential’s quadratic dependence on $u$. The appearance of the absolute values, e.g. $|m+s|$, ensures that $\eta$ is always finite when $u\in\{-1,1\}$, regardless of the values of $m$ and $s$. This is equivalent to enforcing asymptotic boundary conditions on $S(u)$. In Eq. (126c), $-\gamma$ is also a viable solution for $k_{2}$; however, $k_{2}=+\gamma$ is chosen by convention Leaver (1985). The result for $\mathcal{L}{{}_{u}^{\prime}}$ has two schematic forms of relevance,
@@
In Eq. (B), $\text{a}_{0}$ through $\text{a}_{4}$ are constant coefficients defined in Eqs. (129a-129e). They should not to be confused with the similarly labeled coefficients defined in the main text.
@@
Equation (128) defines $\mathcal{D}^{\prime}_{u}$ to be the differential part of Eq. (B) (i.e. Eqs. 127b and 127c). The operator, $\mathcal{D}^{\prime}_{u}$, is useful for understanding the eigenproblem (Eq. 125c) in the non-spinning limit (i.e. $\gamma=a\tilde{\omega}=0$), and for general spins.
@@
In Eqs. (B) and (B), $\text{b}_{0}$ through $\text{b}_{6}$ are $n$ dependent constants defined in Eqs. (143a-143g).
@@
There is no weight function in Eq. (144b) because Eq. (122b) is explicitly in Sturm-Liouville form (with weight function $1$Arfken et al. (2013). The integrands of Eqs. (144a) and (144b) must be equivalent, meaning that $\text{W}^{\text{Jac.}}(u)$ and $\eta(u)|_{\gamma=0}$ are related,
@@
Equations (B-B) can be used to determine the values of $a_{n}$ as follows. Given the scalar product, Eq. (141a), we will use bra-ket notation to simplify to representation of various steps. In that notation, the completeness of the Jacobi polynomials is expressed as
@@
$\hat{L}\;$ =$=\;\{\langle{p_{n^{\prime}}}|\mathcal{L}{{}_{u}^{\prime}}|{p_{n}}\rangle\}_{n,n^{\prime}}\;,$ (150a)
$\vec{g}\;$ =$=\;\{{\langle{p_{n}}\,|\,{g}\rangle}\}_{n}\;.$ (150b)
@@
The essential feature of this approach is that, starting with Eq. (149), the recursion and orthogonality relationships, Eqs. (B-B), can be used to show that $\hat{L}$ is at most tridiagonal. In particular, when $\mathcal{L}{{}_{u}^{\prime}}$ acts on $|{p_{n}}\rangle$, only lower to upper diagonal terms contribute,
@@
For any $N$ dimensional truncation of $\hat{L}$, say $\hat{L}_{N}$, Eq. (B) provides the matrix elements,
@@ Appendix C The confluent Heun equation in natural general form
Perhaps the most referenced (e.g. Refs. Hortacsu (2012); Leaver (1986); Cook (2020)) form of confluent Heun equation is,
@@
[…]
Equation (C) may be derived from the Heun equation (which has $4$ regular singular points) by confluence of two singular points (see e.g. Arfken et al. (2013)) into a single irregular one. Equation (159a) is simply a definition of the confluent Heun equation in terms of its associated linear differential operator, $\mathcal{L}^{(1)}_{z}$. In Eq. (159b), $\alpha$, $\nu$, $\sigma$, $\gamma$ and $\delta$ are complex valued parameters, and $z$ is nominally real valued.
Equation (C), is the confluent Heun equation in what’s called nonsymmetrical canonical form Ronveaux and Arscott (1995); Cook and Zalutskiy (2014); Fiziev (2009b). It happens that this is actually not the most general form of the confluent Heun equation. As a result, a naive comparison between Eq. (C) and Teukolsky’s radial equation may generate unnecessary confusion. To facilitate the upcoming discussion, we will henceforth use $r$ rather than $z$, and we will discuss forms of the confluent Heun equation by referring to the related differential operators, $\mathcal{L}^{(i)}_{\text{cH}}$, such that
@@
=$\mathcal{L}{{}_{r}}\;=\;$ $\left({\text{A}_{0}}+{\text{A}_{1}}r+\text{A}_{2}\,r^{2}+\frac{{\text{A}_{3}}}{r-{r_{-}}}+\frac{{\text{A}_{4}}}{r-{r_{+}}}\right)$ (162)
@@
For completeness, parameters $\text{A}_{1}$ through $\text{A}_{6}$ are defined in terms of $r_{+}$, $r_{-}$ and $s$ as
@@
The fact that Teukolsky’s radial operator is a special case of Heun’s confluent operator is clear by comparing Eq. (161) to Eq. (162). The confluent Heun nature of the operator is not changed under transformations of the dependent or independent variables, but such transformations can impact the physical nature of solutions. For example, application of the QNM boundary conditions (see e.g. Eqs. 11-III.4) has the effect of transforming Eq. (162) into a different version of Eq. (161). However, if the solution of the transformed equation (e.g. Eq. (15)) is to be consistent with QNM boundary conditions, then it cannot have asymptotic behavior that affects a transformation into yet another version of the confluent Heun equation. Application of a compactified coordinate (see Eq. 10) has the effect of changing the coordinate position of singular points, but not their regular or irregular nature.

### v4

@@ VII.2.1 Construction
The polynomial coefficients, $\text{a}_{pkq}$, are determined by applying the form above for $\mathbf{y}_{pk}(\xi)$ to the eigen-relationship given by Eq. (73a). One then asserts that the coefficients of each power in $\xi$ must be equal on both sides of Eq. (73a). The result is a three term recursion relationship or, equivalently, a discrete eigenvalue relationship. To simplify our notation, we will henceforth drop two indices, $p$ and $k$, from $\text{a}_{pkq}$,
@@
Furthermore, if $\mu_{p}$ is given by Eq. (80), then the discrete eigenvalue relationship given by Eq. (76) defines a finite dimensional matrix eigenvalue problem [^17] Note that, along with the series’ boundary conditions (Eqs. 78-79), it is the finite nature of the system of equations that justifies the eigenproblem interpretation. In contrast, non-polynomial solutions to the radial problem will not satisfy Eq. (79), meaning that e.g. a continued fraction based method, and a related boundary condition at the infinite end of the series are required Leaver (1986, 1985).  Ronveaux and Arscott (1995); DLMF (b); Cook and Zalutskiy (2014). The matrix of interest, $\hat{\mathbf{A}}_{p}$, is asymmetric and tridiagonal,
@@ VII.2.2 Orthogonality between polynomials of like order
$(\lambda_{pk}-\lambda_{pk^{\prime}})\,{\langle{\mathbf{y}_{pk^{\prime}}}\,|\,{\mathbf{y}_{pk}}\rangle}\;$ = $=\;0\;.$ (91)
@@
Figure 5: Distribution of roots for confluent Heun polynomials, $y_{pk}(\xi)$, for three cases, all with s=2s=-2, $a/M=0.7$, and $\ell=m=2$. The three cases differ in polynomial order, $p$, and the frequency, $\tilde{\omega}$, used: (left) $y_{33}$, the $n=0$ QNM frequency, $M\tilde{\omega}=0.5326-0.0808i$, was used, (center) $y_{77}$, the $n=0$ QNM frequency, $M\tilde{\omega}=0.4713-0.5843i$, was used, and (right) $y_{20,20}$, the $n=12$ QNM frequency, $M\tilde{\omega}=0.4155-2.5050i$, was used. Polynomial roots are marked by white open circles. Roots were calculated using numpy.roots Horn and Johnson (1990); Harris et al. (2020). Values of polynomial phase, $\arg(y_{pk})$, are shaded from blue to yellow, indicating values between $\pi$ and $\pi$, respectively. Dotted horizontal and vertical lines mark the respective locations of the real and imaginary axes.
@@ VII.2.3 A scalar product expression for eigenvalues
If $y_{pk}$ are known, then Eq. (99) allows the computation of the related eigenvalues via the scalar product. In most instances this will not be needed to define the eigenvalues; however, Eqs. (95-99) are presented to provide a way to check consistency between evaluations of the scalar product, and the underlying confluent Heun system.
@@ VII.2.4 Numerical examples
Numerical computation of the confluent Heun polynomials may proceed by first computing $\text{C}_{2}$, $\text{C}_{3}$ and $\text{C}_{4}$. Then, for a chosen polynomial order, one would compute $\hat{\mathbf{A}}_{p}$ (Eq. 81), so that its eigenvalues and vectors may be determined. Since scalar products between polynomials is a weighted sum over monomial moments (Eq. 64), it is useful to precompute and store all of the monomial moments relevant to a given polynomial order. This allows all polynomial scalar products to be evaluated without explicit integration (Eq. 47). From this perspective, numerical evaluation of confluent Heun polynomials may be accomplished on-the-fly via the appropriate sum over monomials (Eq. 74). Here, this procedure is used to provide select numerical examples.
@@
This effect has two causes, one technical and the other mathematical. The technical cause of this feature is simply the value of the precision parameter, dps=64=64, which means that numerical results near $10^{-64}$ are more prone to round-off error. The mathematical cause of this feature is that, like the legendre polynomials, the confluent Heun polynomials have normalization constants that generally decrease as polynomial order increases [^18] For the Legendre polynomials, the analog of $\eta_{pk}^{2}$ (see Eq. 92) is $(2p+1)^{-1}$.. Consequently the action of normalization effectively scales each polynomial by a very large number, exacerbating finite precision errors. At worst, square normalization constants might be $10^{-64}$ or smaller and would thereby be corrupted by finite precision error. In turn, this would correspond to normalization constants $10^{-32}$ or smaller, and the act of normalization would be to scale the non-normalized polynomials by a corrupted number of order $10^{32}$. This situation can be avoided by increasing numerical precision.
@@
At focus in the left panel of Fig. 7 is the distribution of normalization constants. While normalization constants are found to generally decrease with polynomial order (not shown in figure), they are also found to generally decrease as $k$ increases, and they are found to generally increase as the imaginary part $\tilde{\omega}$ increases.
@@ VIII Implications for analytic solutions to Teukolsky’s radial equation
In effect, Eq. (103) fully defines a solution ansatz. This solution ansatz may be used to investigate whether the scalar product and confluent Heun polynomials imply that Teukolsky’s radial equation is, like the angular equation, exactly tridiagonalizable. To this end, we may attempt to represent the radial equation in the basis defined by $\{u_{p}\}_{p=0}^{\infty}$.
@@
Rewriting $\mathcal{L}{{}_{\xi}}$ in terms of $\mathcal{D}_{\xi}$ (see Eq. 101) allows $\mathcal{L}{{}_{\xi}}\,|{u_{p}}\rangle$ to be expanded as
@@
It follows from Eq. (107c) that $\mathcal{L}{{}_{\xi}}\,|{u_{p}}\rangle$ may be rewritten as
@@
Together, Eq. (103d), Eq. (110) and Eq. (113) allow $\mathcal{L}_{\xi}|{u_{p}}\rangle$ to be written as
@@
The implication of Eqs. (114-115) is that, if the canonical polynomials $u_{n}$ can be determined, then Teukolsky’s radial operator is exactly tridiagonalizable.
@@ Appendix A Notation
$\mathcal{L}{{}_{r}}\;$=\; $V_{R}(r)\;+\;\Delta(r)^{-s}\partial_{r}\left[\Delta(r)^{s+1}\partial_{r}\right]\;,$ (119)
@@ Appendix B Solving Teukolsky’s angular equation with Jacobi polynomials
Equivalently, $k_{0}$, $k_{1}$ and $k_{2}$ are the singular exponents that may be found via the method of Frobenius Courant and Hilbert (1989); Arfken et al. (2013); Pinchover and Rubinstein (2005). In effect, $k_{0}$ and $k_{1}$ remove divergences from the potential at the regular singular points, and $k_{2}$ removes the potential’s quadratic dependence on $u$. The appearance of the absolute values, e.g. $|m+s|$, ensures that $\eta$ is always finite when $u\in\{-1,1\}$, regardless of the values of $m$ and $s$. This is equivalent to enforcing asymptotic boundary conditions on $S(u)$. In Eq. (126c), $\gamma$ is also a viable solution for $k_{2}$; however, $k_{2}=+\gamma$ is chosen by convention Leaver (1985). The result for $\mathcal{L}{{}_{u}^{\prime}}$ has two schematic forms of relevance,
@@
In Eq. (127), $\text{a}_{0}$ through $\text{a}_{4}$ are constant coefficients defined in Eqs. (129a-129e). They should not to be confused with the similarly labeled coefficients defined in the main text.
@@
Equation (128) defines $\mathcal{D}^{\prime}_{u}$ to be the differential part of Eq. (127) (i.e. Eqs. 127b and 127c). The operator, $\mathcal{D}^{\prime}_{u}$, is useful for understanding the eigenproblem (Eq. 125c) in the non-spinning limit (i.e. $\gamma=a\tilde{\omega}=0$), and for general spins.
@@
In Eqs. (140) and (141), $\text{b}_{0}$ through $\text{b}_{6}$ are $n$ dependent constants defined in Eqs. (143a-143g).
@@
There is no weight function in Eq. (144b) because Eq. (122b) is explicitly in Sturm-Liouville form (with weight function 11Arfken et al. (2013). The integrands of Eqs. (144a) and (144b) must be equivalent, meaning that $\text{W}^{\text{Jac.}}(u)$ and $\eta(u)|_{\gamma=0}$ are related,
@@
Equations (140-143) can be used to determine the values of $a_{n}$ as follows. Given the scalar product, Eq. (141a), we will use bra-ket notation to simplify to representation of various steps. In that notation, the completeness of the Jacobi polynomials is expressed as
@@
$\hat{L}\;$ $=\;\{\langle{p_{n^{\prime}}}|\mathcal{L}{{}_{u}^{\prime}}|{p_{n}}\rangle\}_{n,n^{\prime}}\;,$ (150a)
$\vec{g}\;$ $=\;\{{\langle{p_{n}}\,|\,{g}\rangle}\}_{n}\;.$ (150b)
@@
The essential feature of this approach is that, starting with Eq. (149), the recursion and orthogonality relationships, Eqs. (140-143), can be used to show that $\hat{L}$ is at most tridiagonal. In particular, when $\mathcal{L}{{}_{u}^{\prime}}$ acts on $|{p_{n}}\rangle$, only lower to upper diagonal terms contribute,
@@
For any $N$ dimensional truncation of $\hat{L}$, say $\hat{L}_{N}$, Eq. (152) provides the matrix elements,
@@ Appendix C The confluent Heun equation in natural general form
Perhaps the most referenced (e.g. Refs. Zarrinkamar and Hortaçsu (2018); Leaver (1986); Cook (2020)) form of confluent Heun equation is,
@@
$\;+\,$ $\left[\gamma(z-1)+z(\delta+\nu(z-1))\right]\,\partial_{z}$
$\;+\,$ $(z-1)z\,\partial^{2}_{z}\;.$
Equation (159) may be derived from the Heun equation (which has $4$ regular singular points) by confluence of two singular points (see e.g. Arfken et al. (2013)) into a single irregular one. Equation (159a) is simply a definition of the confluent Heun equation in terms of its associated linear differential operator, $\mathcal{L}^{(1)}_{z}$. In Eq. (159b), $\alpha$, $\nu$, $\sigma$, $\gamma$ and $\delta$ are complex valued parameters, and $z$ is nominally real valued.
Equation (159), is the confluent Heun equation in what’s called nonsymmetrical canonical form Ronveaux and Arscott (1995); Cook and Zalutskiy (2014); Fiziev (2009b). It happens that this is actually not the most general form of the confluent Heun equation. As a result, a naive comparison between Eq. (159) and Teukolsky’s radial equation may generate unnecessary confusion. To facilitate the upcoming discussion, we will henceforth use $r$ rather than $z$, and we will discuss forms of the confluent Heun equation by referring to the related differential operators, $\mathcal{L}^{(i)}_{\text{cH}}$, such that
@@
$\mathcal{L}{{}_{r}}\;$=\; $\left({\text{A}_{0}}+{\text{A}_{1}}r+\text{A}_{2}\,r^{2}+\frac{{\text{A}_{3}}}{r-{r_{-}}}+\frac{{\text{A}_{4}}}{r-{r_{+}}}\right)$ (162)
@@
For completeness, parameters $\text{A}_{1}$ through $\text{A}_{6}$ are defined in terms of $_{+}, rr_{-}$ and $s$ as
@@
The fact that Teukolsky’s radial operator is a special case of Heun’s confluent operator is clear by comparing Eq. (161) to Eq. (162). The confluent Heun nature of the operator is not changed under transformations of the dependent or independent variables, but such transformations can impact the physical nature of solutions. For example, application of the QNM boundary conditions (see e.g. Eqs. 11-13) has the effect of transforming Eq. (162) into a different version of Eq. (161). However, if the solution of the transformed equation (e.g. Eq. (15)) is to be consistent with QNM boundary conditions, then it cannot have asymptotic behavior that affects a transformation into yet another version of the confluent Heun equation. Application of a compactified coordinate (see Eq. 10) has the effect of changing the coordinate position of singular points, but not their regular or irregular nature.